# The time scale of redundancy between prosody and linguistic context

### v2

@@ 5 Results
Fig. 1 displays the average mutual information (MI) across the 6 prosodic features tested here (pitch, ludness, duration, absolute prominence, relative prominence, pause), and Fig. 3 displays those results for each feature separately. See App. A for the values of unconditional entropies that comprise those MIs. Our results’ overall trend confirms Hypothesis 1, the long-scale past redundancy hypothesis. When averaged across all linguistic features, the mutual information between prosody and past linguistic context (first column in Fig. 1) increases as a function of the number of words available, up to about 5-8 words, after which it plateaus. Our results also support Hypothesis 2, the short-scale future redundancy hypothesis. When averaged across all linguistic features, the mutual information between prosody and future linguistic context (first row in Fig. 1) increases only up to about one or two words and then plateaus. Furthermore, the MI with the past is higher than the MI with the future, a trend that becomes larger for longer spans of words.
@@
When examining the mutual information with linguistic contexts containing both past and future words (i.e., $n>0$ and $m>0$; the $n$-th column and $m$-th row in Fig. 1), we observed interesting interactions. Specifically, a combination of about one word in the future and about 5-8 words in the past already contains most of the information about prosodic features. Notably, this $n,m$ combination led to higher mutual information than even other combinations with larger context (i.e., with $n^{\prime},m^{\prime}$, $n^{\prime}\!\geq\!n$, and $m^{\prime}\!\geq\!m$). While this is theoretically impossible (adding more context can never decrease mutual information), this is likely due to our models’ training procedure not being able to ignore unhelpful contributions of long-scale contexts.
@@
Interestingly, each individual prosodic feature shows a slightly different pattern (see Fig. 3). The long-scale past hypothesis (Hypothesis 1) is supported for most prosodic features individually, but not for duration and pauses; for prominence, pitch and loudness, the MI with past context (red curves in the lower plot per each feature in Fig. 3) increases up to at least $3$ words.111111 For absolute prominence, the curve is a little noisy but saturates at around 7 words. For pauses though, the past MI saturates after only two words, and for duration, the MI with past alone does not seem to significantly rise above the $0,0$ point, indicating the past context does not add information beyond the word identity itself. However, for all prosodic features, even duration, when looking at all $n,m$ combinations, redundancy is largest around 1-2 words in the future and 3-8 words in the past—thus supporting Hypothesis 1.
@@
The short-scale future hypothesis (Hypothesis 2) is also supported for most individual features, except for duration and pause. For prominence, pitch and loudness, the time scale of the future MI is shorter than the scale of the past MI, saturating somewhere between 1 and 3 words (blue curves in the lower plot per each feature in Fig. 3). For pauses though, the future MI saturates after about 4 words, which is longer than for the past—although these MI estimates seem noisy. For duration, it saturates after 1 word, which is more than the past since the past curve shows no increasing trend. Notably however, the MI with the past is larger than with the future in all features except for duration, and, even for duration, ${\mathrm{MI}}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}},{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}_{\overset{n,1}{\leftrightarrow}}}})$ is larger than ${\mathrm{MI}}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{P}_{t}}},{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{W}}}}}}_{\overset{1,m}{\leftrightarrow}}}})$. We thus conclude that all these individual feature results’ support Hypothesis 2.
@@ 6 Discussion
This work aimed to estimate the time scale of the redundancy between prosodic and linguistic information. We built on an existing published pipeline Wolf et al. (2023), but extended it to include a systematic modulation of context length, from 1 to 10 words. We will make our extension of this pipeline publicly available for researchers wishing to extend our analysis, potentially exploring the time scale of redundancy between other communication channels of interest.
Overall, we confirm our two hypotheses (see Section § 1): For most prosodic features we tested (apart from pause and duration, see § 6.4), redundancy with past linguistic context unfolds across a long time scale (of about 3-8 words), while redundancy with future words is shorter-scale (concentrated on 1-2 future words). We next discuss the implications of these results with respect to previous literature.
6.1 The Time Scale of Prosodic Information
Sentence comprehension is constrained by cognitive demands such as attention and working memory, leading listeners to maintain a lossy representation of past linguistic context Gibson (1998); Vasishth et al. (2010); Futrell et al. (2020); Kuribayashi et al. (2022). While the precise number of words maintained in working memory probably depends on many factors, estimates suggest a range of about 3-5 words Cowan (2010). In the brain, language-selective neural populations integrate linguistic information across distinct time scales; these scales are quantified to span between 1-6 words (Jain and Huth 2018; Regev et al. 2024).
Despite the well-studied dynamics of linguistic processing, the time scale at which prosody interacts with linguistic information had been previously underexplored. Intonation units, a meaningful organizational unit of prosodic information, follow a rhythmic structure of about 1 Hz Inbar et al. (2020); meaning that each unit is about 1 second long and therefore contains about 3-4 words, given an average speech rate of about 200 words per minute Yuan et al. (2006). This, together with the span of working memory (mentioned above), suggests a natural alignment between prosodic structure and the cognitive constraints of linguistic processing.
Here, we provide a quantification of the time scale at which redundancy between prosody and linguistic context operates. Our findings suggest that this redundancy spans around 3-8 past words, a scale comparable to both linguistic working memory limitations and prosodic segmentation. These results thus highlight a possible role of prosody in optimizing comprehension.
@@ 6.2 Prosody as an Audience-Design Tool in Communication
A longstanding debate in linguistics concerns the extent to which language production is shaped by audience design, with speakers actively tailoring their utterances to facilitate listener comprehension. Evidence suggests that syntactic choices are not strongly adapted for listener needs but rather reflect the speaker’s own constraints Ferreira (2008); Morgan and Ferreira (2022). This apparent lack of audience design in syntax may stem from the rigid structural constraints imposed by linguistic systems. In contrast, prosody may offer greater flexibility, allowing speakers to dynamically modulate pitch, loudness, and rhythm in real time. This flexibility suggests that prosody may play a larger role in audience design Clark et al. (2025), serving as an additional communicative channel that enhances intelligibility. In fact, prior work shows a trade-off between a word’s duration and its information content (Jurafsky et al., 1998; Bell et al., 2009; Coupé et al., 2019; Pimentel et al., 2021), a trade-off which is typically interpreted as arising to facilitate listeners comprehension by smoothing the amount of information they receive per second (known as the uniform information hypothesis; Fenk and Fenk, 1980; Genzel and Charniak, 2002; Levy and Jaeger, 2007). Our findings support this audience-design view by revealing that the redundancy between prosody and past linguistic context extends over long time scales, suggesting that prosody may serve as a "reminder", helping listeners access information from the long-scale past.
@@ 6.3 The Relationship Between Prosody and Future Words
Our findings indicate that redundancy between prosody and linguistic information is weaker for future words than for past words. However, prosody still exhibits a strong relationship with the immediately upcoming word or two. This short-range relationship could stem from motor constraints on prosody production, or from local linguistic dependencies like fixed expressions. Another possibility is that prosodic planning occurs at the level of entire sentences and therefore observed redundancy with the next word reflects broader contour structuring. Finally, prosody may actively signal upcoming words through cues such as duration, pauses, or pitch changes, aiding listener expectations. Future work should explore these potential mechanisms to better understand prosody’s role in forward-looking processing. Notably, prior research has shown that a word’s duration correlates with its predictability given future context (Bell et al., 2009). Further, even reading times—a setting in which the subject is assumed to not know what future words are—correlate with features of future words (such as frequency, predictability and entropy; Roark et al., 2009; Angele et al., 2015; van Schijndel and Schuler, 2017).
6.4 Pause and Duration Do not Follow Hypotheses 1 and 2
Compared to the other prosodic features, pause and duration stand out in their relatively short time scale of redundancy with both past and future words, as well as a relatively strong redundancy with future words. This may suggest that pausing after a word and elongating it, are mainly served in order to prepare for the next upcoming word and facilitate its processing by slowing down the rhythm of speech. Alternatively, as discussed above, pauses are most common at the end of sentences. Similarly, sentence-final words are also typically elongated Seifart et al. (2021); Paschen et al. (2022). The high predictability of both these features with future context could be partially due to our models predicting a sentence-final vs sentence-middle distinction, which could itself be used to roughly predict the value of these prosodic features.
@@ 7 Conclusion
Our findings reveal a fundamental asymmetry in the time scale of redundancy between prosody and linguistic context: while prosody exhibits redundancy with both past and future words, this relationship extends across a longer span for past words (3–8 words) than for future words (1–2 words). This suggests that prosody’s relationship with future words primarily reflects short-term effects such as next-word prediction, local word dependencies, or other production factors—future work should try to distinguish those explanations. In contrast, its relationship with past words operates over a broader scale, potentially serving to reinforce or highlight information that may be cognitively demanding for listeners to remember in real-time communication. These results provide new insights into the role of prosody in spoken language.
@@
The first set of limitations relates to the dataset used. Our dataset consists of audiobooks, which do not necessarily reflect natural prosody in real time communication, potentially affecting the generalizability of our findings. Redundancy may be higher in audiobooks than in spontaneous speech, because the text is written with the assumption that it must convey all necessary information without relying on prosody. We address this concern to some extent by removing punctuation marks, which serve as a substitute for prosody in written text. Another dataset-related limitation is the sample size. Larger datasets may be required for more stable estimations, especially given that we compute 55 different values (for $\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}_{\overset{n,m}{\leftrightarrow}}}$, $n$ and $m$ from 1 to 10), each based on different subsets of the data, effectively reducing the number of samples available for each estimate. This sparsity likely contributes to some of the observed noise in our results.
@@
The second set of limitations has to do with the estimation procedure. The mutual information we compute approximates the true value, and is constrained by the quality of the models we use $p_{\boldsymbol{\theta}}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{p}_{t}}}\mid{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}_{\overset{n,m}{\leftrightarrow}}}})$. One of our modeling assumptions is the functional form of the conditional distribution of prosody given a linguistic context (namely, Gaussian, Gamma or Laplace distributions depending on the prosodic feature). However, this parametric assumption may limit the model’s performance and future work should explore alternative conditional distributions which may improve results. This assumption is particularly violated for features that manifest different distributions; pause, for instance, takes a 0 value in $89.4\%$ of the data and may therefore be better modeled by a zero-inflated distribution. Indeed, our results for pause seem particularly noisy. Additionally, to estimate $\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}_{\overset{n,m}{\leftrightarrow}}}$, we provided the models with short segments of 1 to 10 words. However, these large language models were not pretrained on such short segments but rather on longer spans of text, which might have impacted their efficiency in extracting the information from short segments. While finetuning likely helped mitigate this issue, it remains a potential limitation. Furthermore, we train a single model for all combinations of $n,m$, which does not guarantee that the value is optimal for each combination separately. Finally, we observed cases where the mutual information decayed for longer contexts, which contradicts expectations from information theory, as additional context can never reduce information. This phenomenon likely stems from issues in training the models, which could be biased toward under-utilizing the available context for longer spans. Future work should address these limitations to refine our understanding of redundancy between prosody and linguistic information.

### v3

@@ 5 Results
We present our main results in Figs. 1 and 3. Namely, Fig. 1 displays the average mutual information (MI) across the 6 prosodic features tested here (pitch, loudness, duration, pause, absolute prominence, relative prominence), while Fig. 3 displays those results for each feature separately. See section App. A for the values of the unconditional entropies comprising those MIs.
The overall trend is consistent with the hypotheses outlines above. In line with the long-scale past redundancy hypothesis (Hypothesis 1), the mutual information between prosody (averaging across all prosodic features used here) and past context (first column in Fig. 1) increases as a function of context length until up to about 5-8 words, at which point it plateaus. In line with the short-scale future redundancy hypothesis (Hypothesis 2), the $\mathrm{MI}$ between prosody (averaging across all features) and future context (first row in Fig. 1) increases only until up to one or two words and then plateaus. Furthermore, at context lengths above one word, the MI with the past context tends to be higher than the MI with the future context, with this trend becoming more pronounced at longer contexts.
@@
We also empirically explore other combination of past and future contexts (i.e., $n>0$ and $m>0$; the $n$-th column and $m$-th row in Fig. 1). This analysis revealed that prosodic features at the current word are best predicted by a combination of 5-8 words in the past and about one word in the future. Notably, this $n,m$ combination led to higher mutual information than even other combinations with larger context (i.e., with $n^{\prime},m^{\prime}$, $n^{\prime}\!\geq\!n$, and $m^{\prime}\!\geq\!m$). While this is theoretically impossible (adding more context can never decrease mutual information), this is likely due to our models’ training procedure not being able to ignore unhelpful contributions of long-scale contexts.
@@
Interestingly, individual prosodic features show somewhat distinct patterns (see Fig. 3). The long-scale past hypothesis (Hypothesis 1) is supported for most prosodic features individually, but not for duration and pauses. For pitch, loudness, and prominence, the MI with past context (red curves in the lower plot for each feature in Fig. 3) increases until up to at least $3$ words (although the curve for absolute prominence is a little noisy). For duration, however, the MI with past context never rises significantly above the $0,0$ point, indicating the past context does not add information beyond the word identity itself; and for pauses, the past MI saturates after two words.
@@
The short-scale future hypothesis (Hypothesis 2) is also supported for most individual prosodic features, except for duration and pause. For pitch, loudness, and prominence, the MI with future context is shorter than with past context, saturating between 1 and 3 words (blue curves in the lower plot in each feature in Fig. 3). For duration, the MI with future context saturates after 1 word, which is longer than with past context (the past curve shows no increasing trend). For pauses, the future MI saturates after about 4 words, which is longer than with past context—although these MI estimates are noisy for this feature, likely due to data sparsity. Notably however, the MI with past context is higher than with future context for all prosodic features except for duration. We thus conclude that the results from these individual prosodic features generally support the short-scale future hypothesis.
@@ 6 Discussion
This work aimed to characterize the time scale of the redundancy between prosodic and linguistic information. We built on the foundation of past work, which established that linguistic information is partially redundant with prosodic features Wolf et al. (2023), to systematically manipulate the length of past and future context, from 1 to 10 words. Our extension of this pipeline is publicly available for researchers wishing to extend our analysis, potentially exploring the time scale of redundancy between other communication channels of interest.
Based on what is known about human memory limitations and linguistic planning, we hypothesized that redundancy between linguistic information and prosodic features extends over a relatively long time scale into the past, and a relatively short time scale into the future (see section § 1). We find support for these hypotheses: For most prosodic features we tested (apart from duration and pause, see section § 6.4), redundancy with past linguistic context unfolds across a long time scale (of 3-8 words), whereas redundancy with future linguistic context is shorter-scale (of 1-2 words). We next discuss the implications of these results with respect to the previous literature.
6.1 The Time Scales of Linguistic and Prosodic Information
Sentence comprehension is constrained by cognitive demands such as attention and working memory, leading to lossy representations of past linguistic context (Gibson, 1998; Vasishth et al., 2010; Futrell et al., 2020; Kuribayashi et al., 2022). It is, however, challenging to estimate the precise number of words maintained in working memory during language processing, as this likely depends on many factors. Neural evidence suggests that language-selective neural populations [^6] Language-selective neural populations are clusters of neurons in the human brain that respond selectively to linguistic experimental manipulations, but not to manipulations in other perceptual or cognitive domains such as math and music. These populations are thought to constitute a brain network specialized for language processing Fedorenko et al. (2024). integrate linguistic information across time scales varying between 1-6 words (Jain and Huth, 2018; Shain, Kean, et al., 2024; Regev, Casto, et al., 2024). Moreover, work in prosody has put forward the notion of "intonation units", which follow a rhythmic structure of about 1 Hz (Inbar et al., 2020). Given the average speech rate of  200 words per minute (Yuan et al., 2006), each intonation unit contains about 3-4 words, which aligns with the relatively short context of linguistic processing.777 To be precise about the generality of the claims above, we note that: the claims in (Gibson, 1998; Vasishth et al., 2010; Futrell et al., 2020; Kuribayashi et al., 2022) are intended as universal and were shown for English, German, and Japanese; the experiments in (Jain and Huth, 2018; Shain, Kean, et al., 2024; Regev, Casto, et al., 2024) are in English; the experiments in (Inbar et al., 2020) cover 6 typologically diverse languages; the experiments in (Yuan et al., 2006) are in English and Chinese.
Our findings are generally in accord with these previous studies, and suggest that redundancy with past context spans around 3-8 words—a scale comparable to both the time scale of linguistic integration and prosodic segmentation.
@@ 6.2 Prosody as an Audience-Design Tool in Communication
A longstanding debate in linguistics concerns the extent to which language production is shaped by audience design considerations, with speakers tailoring their utterances to facilitate comprehension. Some evidence suggests that syntactic choices are not strongly adapted for listener needs but rather reflect the speaker’s own constraints Ferreira (2008); Morgan and Ferreira (2022). This apparent lack of audience design in syntax may stem from the rigid structural constraints imposed by the rules of the language (at least in English, which is used in much past work). In contrast, prosodic choices may offer greater flexibility for accommodating audience-centered considerations, allowing speakers to dynamically modulate pitch, loudness, and rhythm in real time. This flexibility may afford prosody a larger role in audience design Clark et al. (2025), serving as a communicative channel that facilitates comprehension. For example, prior work shows a trade-off between a word’s duration and its information content (Jurafsky et al., 1998; Bell et al., 2009; Coupé, Oh, et al., 2019; Pimentel et al., 2021), which is typically interpreted as arising to facilitate comprehension by more evenly distributing information across time (known as the uniform information density hypothesis; Fenk and Fenk, 1980; Genzel and Charniak, 2002; Levy and Jaeger, 2007; Meister et al., 2021). Our findings may be taken as further evidence for the audience-design view of prosody: the fact that its redundancy with past context extends over long time scales suggests that prosody may help clarify the relationship between the current word and the broader preceding context, thus facilitating the extraction of the intended meaning. Given that prosody appears to carry locally unique information, prosodic features may aid comprehension when representations of longer-scale past are lossy.
@@ 6.3 The Relationship Between Prosody and Future Words
Our findings indicate that the redundancy between prosody and future words is overall lower compared to past words. However, prosody still exhibits a strong relationship with the upcoming word or two. This short-range relationship may arise from motor constraints on prosodic articulation, from local linguistic dependencies such as fixed expressions, or from short-scale prosodic planning—consistent with theories of language production that emphasize incremental, short-term planning Brown-Schmidt and Konopka (2008); Bock et al. (1994). In addition, prosody may help listeners form expectations about upcoming words through cues such as duration, pauses, or pitch modulations. This interpretation aligns with recent findings that humans are better at predicting upcoming words given spoken, compared to written, context Botch and Finn (2025). Relatedly, prior research has shown that a word’s duration correlates with its predictability given future context (Bell et al., 2009). Along similar lines, reading times also seem to correlate with features of upcoming words (such as frequency, predictability and entropy; Roark et al., 2009; Angele et al., 2015; van Schijndel and Schuler, 2017). Future work should investigate these potential mechanisms to clarify the role of prosody in forward-looking processing.
6.4 Duration and Pause Show Weak Past, Strong Future Redundancy
Compared to the other prosodic features, duration and pause stand out in their relatively short time scale of redundancy with both past and future words, as well as in their relatively high MI with future words. This pattern may suggest that lengthening a word and pausing after it primarily serves to prepare for the upcoming word, facilitating its processing by slowing the rhythm of speech. Alternatively, this pattern may stem from sentence boundaries: As discussed above, pauses are most common after sentence-final words, which are also often elongated Seifart et al. (2021); Paschen et al. (2022). The high predictability of duration and pause from future context may therefore partly reflect our models’ ability to detect a sentence-final vs. sentence-medial word, an important structural cue which could, in turn, help predict the values of these prosodic features. While we removed punctuation from the input, therefore preventing sentence-boundary information from being conveyed explicitly through punctuation marks, we did not lowercase all the data and therefore sentence-initial words are clearly detectable by their capitalization. Furthermore, sentence boundary cues could be inferred from the lexical content itself, since certain words tend to occur more frequently at the beginning or end of sentences.
@@ 7 Conclusion
Our findings reveal a fundamental asymmetry in the time scale of redundancy between prosody and linguistic information: while prosody exhibits redundancy with both past and future words, this relationship extends across a longer span for past words (3–8 words) than for future words (1–2 words). This suggests that prosody’s relationship with future words primarily reflects short-term effects such as next-word prediction, local word dependencies, or other production factors—future work should try to distinguish those explanations. In contrast, its relationship with past words operates over a broader scale, potentially serving to reinforce or highlight information that may be cognitively demanding for listeners to process in real-time communication. These results provide new insights into the role of prosody in spoken language.
@@
The first set of limitations relates to the dataset used. Our dataset consists of audiobooks, which do not necessarily reflect natural prosody in real time communication, potentially affecting the generalizability of our findings. Redundancy may be higher in audiobooks than in spontaneous speech, because the text is written with the assumption that it must convey all necessary information without relying on prosody. We address this concern to some extent by removing punctuation marks, which serve as a substitute for prosody in written text. Another dataset-related limitation is the sample size. Furthermore, the syntactic structures used in audiobooks may differ from those in spontaneous speech. Because pause and duration often signal syntactic boundaries, they may be particularly sensitive to such structural differences. This may help explain why pause and duration did not follow our hypotheses. Larger datasets may be required for more stable estimations, especially given that we compute 55 different values (for $\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}_{\overset{n,m}{\leftrightarrow}}}$, $n$ and $m$ from 0 to 9). Furthermore, our dataset is limited to the English language. Therefore, our findings are English-specific and may not generalize across languages, given known cross-linguistic variation in prosody. Future work should generalize this work to other typologically distinct languages (see Wilcox et al., 2025, for work using similar methods to study diversity in prosodic typology).
@@
The second set of limitations has to do with the estimation procedure. The mutual information we compute tries to approximate the true value, and is constrained by the quality of our models $p_{\boldsymbol{\theta}}({{\color[rgb]{0,0.4765625,0.80078125}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4765625,0.80078125}\mathrm{p}_{t}}}\mid{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}_{\overset{n,m}{\leftrightarrow}}}})$. One of our modeling assumptions is the functional form of the conditional distribution of prosody given a context (namely, Gaussian, Gamma or Laplace distributions depending on the prosodic feature). However, this parametric assumption may limit the model’s performance and future work should explore alternative conditional distributions which may improve results (as done by, e.g., Wilcox et al., 2025). This assumption is particularly violated for features that manifest different distributions; pause, for instance, takes a 0 value in $89.4\%$ of the data and may therefore be better modeled by a zero-inflated distribution. Indeed, our results for pause seem particularly noisy. Additionally, to estimate $\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\boldsymbol{{{\color[rgb]{0.2734375,0.62890625,0.28515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.62890625,0.28515625}\mathrm{w}}}}}}_{\overset{n,m}{\leftrightarrow}}}$, we provided the models with short segments of 1 to 10 words. However, the language models used here (e.g., BERT) were not pretrained on such short segments, but rather on longer spans of text; this might have impacted their efficiency in extracting the information from short segments. While finetuning likely mitigates this issue, it remains a potential limitation. Furthermore, we train a single model for all combinations of $n,m$, which does not guarantee that the model is optimal for each combination separately. Finally, we observed cases where the mutual information decayed for longer contexts, which contradicts expectations from information theory, as additional context can never reduce information. This phenomenon likely stems from issues in training the models, which could be biased toward under-utilizing the available context for longer spans. Future work should address these limitations to refine our understanding of redundancy between prosody and linguistic information.